In a geometric sequence, the second term exceeds the fourth term by 1/2 .
The third term exceeds the fifth term by 1/8. Find the sum to infinity.
1
Expert's answer
2020-12-07T09:13:13-0500

Let "a" and "r" be the first term and common ratio of the geometric sequence.

We know that n-th term of the geometric sequence is

"Tn=a.r^(n-1)"

So,"T2=a.r,T3=a.r^2,T4=a.r^3,T5=a.r^4"

Now given that,"T2-T4=1\/2.........(1)" &

"T3-T5=1\/8........(2)"

From (1) , "ar-ar^3=1\/2"

=> "ar(1-r^2)=1\/2.......(3)"

From (2), "ar^2-ar^4=1\/8"

=> "ar^2(1-r^2)=1\/8........(4)"

Now divide (4) by (3),we get

"ar^2(1-r^2)\/ar(1-r^2)=(1\/8)\/(1\/2)"

=> "r=1\/4"

From (3),putting the value of "r" we get,

"a(1\/4-1\/64)=1\/2"

=> "a.(15\/64)=1\/2"

=> "a=32\/15"

Now we know that sum of the infinite geometric sequence is "=a\/(1-r)" .

Therefore sum of the required geometric sequence is "=(32\/15)\/(1-(1\/4))"

"=(32\/15)\/(3\/4)"

"=(32\/15)\u00d7(4\/3)"

"=128\/45"